# Canonical Autoregressive Generation

### v1

@@ Appendix A Proof of Theorem 1
    The first $i$ merges in $M_{\mathbf{s}\shortmid t}$ are not the same as in $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$. This means that there exists at least one merge $m\in M_{\mathbf{s}\shortmid t}$ among the first $i$ merges in $M_{\mathbf{s}\shortmid t}$ such that $m\notin[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$. For any such merge $m$ as $m\notin[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, it must hold that $m\in[M_{\mathbf{s}\shortmid t}]_{\bm{t}}$ by Lemma 1. So, in $M_{\mathbf{s}\shortmid t}$, merge $m_{i}$ is preceded by the first $i-1$ merges of $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and merge $m$. By Lemma 1, $m$ does not affect the tokens that will create $\mathbf{s}$, so the only merges in $M_{\mathbf{s}\shortmid t}$ before $m_{i}$ that affect $\mathbf{s}$ are the first $i-1$ merges of $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, which are the same as $M_{\mathbf{s}^{\prime}}$. Then, as in case 1, merge $m_{i}$ being applied instead of $m^{\prime}_{i}\prec m_{i}$, implies that $M_{\mathbf{s}\shortmid t}$ cannot be the merge sequence that creates the canonical tokenization of $\texttt{dec}(\mathbf{s}\shortmid t)$ according to BPE.
@@
We will now examine the case where $m_{i}=m^{\prime}_{i}$ for all $i\in[\min(n,n^{\prime})]$, $m_{i}\in[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, $m^{\prime}_{i}\in M_{\mathbf{s}^{\prime}}$ but $n\neq n^{\prime}$. If $n>n^{\prime}$, then there exists at least one merge that can be applied on $\mathbf{s}^{\prime}$ after all merges of $M_{\mathbf{s}^{\prime}}$ are done, which means that $\mathbf{s}^{\prime}$ cannot be canonical. If $n^{\prime}>n$, then there exists at least one merge that can be applied on $\mathbf{s}$ after all merges of $[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ are done. This merge can also be applied on $\mathbf{s}\shortmid t$, which means that $\mathbf{s}\shortmid t$ cannot be canonical.
@@
All cases lead to a contradiction, which implies that $\mathbf{s}\shortmid t$ is non-canonical. ∎
Appendix B Unigram- and Wordpiece-based Tokenizers
Here, we review the Unigram [4] and Wordpiece [5] tokenization algorithms, and prove that appending any token to a non-canonical token sequence results in a sequence that is also non-canonical, according to tokenizers based on these algorithms.
B.1 The Unigram tokenization algorithm
After the initial vocabulary has been set, the algorithm proceeds in iterations, each time computing a loss over the strings in $\mathcal{C}$ and the current vocabulary, and removing a batch of tokens from the vocabulary (typically $10\%$ or $20\%$ of tokens) whose removal minimizes this loss. In each iteration, every token $t$ in the current vocabulary $V$ is assigned a probability score $r(t)=\frac{freq(t)}{\sum_{t^{\prime}\in V}freq(t^{\prime})}$, where $freq(t)$ denotes the number of times that the token $t$ appears in all possible tokenizations of the strings in $\mathcal{C}$. For each token $t\in V$, the loss over the training set is computed as $\sum_{\bm{\sigma}\in\mathcal{C}}-\log(r_{V\setminus\{t\}}(\bm{\sigma}))$, where $r_{V}(\bm{\sigma})=\max_{\mathbf{s}\in V^{*},\texttt{dec}(\mathbf{s})=\bm{\sigma}}r(\mathbf{s})$ denotes the probability score of the most likely tokenization of $\bm{\sigma}$ under vocabulary $V$, and the probability score of tokenization $\mathbf{s}=t_{1}\shortmid\dots\shortmid t_{n}$, with $n\in\mathbb{N}$, is simply $r(\mathbf{s})=r(t_{1})\dots r(t_{n})$. The tokens that minimize this loss are removed from the vocabulary and the process repeats until the vocabulary reaches a predetermined size.
After the vocabulary has been finalized, the encoder is set to tokenize a string $\bm{\sigma}\in\Sigma^{*}$ by finding its most likely tokenization under the final vocabulary $V$, i.e., $\texttt{enc}(\bm{\sigma})=\arg\max_{\mathbf{s}\in V^{*},\texttt{dec}(\mathbf{s})=\bm{\sigma}}r(\mathbf{s})$, using the Viterbi algorithm [38], and the decoder decodes all tokens in $V$ the same way as in the original, large vocabulary.
@@ B.1 The Unigram tokenization algorithm
Let ${\mathcal{T}}=(\Sigma,V,\texttt{enc},\texttt{dec})$ be a Unigram-based tokenizer and $\mathbf{s}\in V^{*}$ a non-canonical token sequence according to ${\mathcal{T}}$. Then, for any token $t\in V$, $\mathbf{s}\shortmid t$ is also non-canonical.
@@
If $\mathbf{s}$ is non-canonical according to Unigram, then let $\mathbf{s}^{\prime}$ denote the canonical tokenization of the same character string, $\texttt{dec}(\mathbf{s})=\texttt{dec}(\mathbf{s}^{\prime})$. Because $\mathbf{s}^{\prime}$ is canonical, it must be that $r(\mathbf{s}^{\prime})>r(\mathbf{s})$. It follows that $r(\mathbf{s}\shortmid t)=r(\mathbf{s})r(t)<r(\mathbf{s}^{\prime})r(t)=r(\mathbf{s}^{\prime}\shortmid t)$, so $\mathbf{s}\shortmid t$ cannot be the canonical tokenization of $\texttt{dec}(\mathbf{s}\shortmid t)$. ∎
B.2 The Wordpiece tokenization algorithm
In the initialization phase, $\Sigma$ is set to contain all characters that appear at least once in the training set of strings $\mathcal{C}$. Then, for each character $c\in\Sigma$ that appears at least once in $\mathcal{C}$, a single-character token $t$ is added to $V$ such that $\texttt{dec}(t)=c$, and ${\mathcal{S}}$ is initialized as a set that contains all single-character token sequences $\mathbf{s}\in V^{*}$ that represent all strings in $\mathcal{C}$. Interestingly, Wordpiece transforms characters (and substrings) inside words differently than characters (and substrings) at the beginning of words. Specifically, tokens representing characters (and substrings) inside words have a special prefix.
To build the vocabulary, Wordpiece proceeds iteratively by merging existing tokens and adding them to $V$ until it reaches a predetermined size, similarly to BPE. However, the criterion to select which pair of tokens to merge is different. If $freq(\mathbf{s}^{\prime})$ denotes the number of times that sequence $\mathbf{s}^{\prime}\in V^{*}$ appears (as a subsequence) in the set of sequences ${\mathcal{S}}$, Wordpiece looks for the pair of tokens $t,t^{\prime}\in V$ that maximizes the value of $\frac{freq(t\shortmid t^{\prime})}{freq(t)\cdot freq(t^{\prime})}$. Then, a new token $t\circ t^{\prime}$ is added to $V$, where $\texttt{dec}(t\circ t^{\prime})=\texttt{dec}(t)\shortmid\texttt{dec}(t^{\prime})$, and all occurences of $t\shortmid t^{\prime}$ in each token sequence $\mathbf{s}\in{\mathcal{S}}$ are replaced by $t\circ t^{\prime}$. With this criterion, Wordpiece prefers to merge tokens whose concatenation appears commonly in ${\mathcal{S}}$, but they are not common individually.
After the above iterative process terminates, the algorithm defines the encoder and decoder functions as follows. For any token sequence $\mathbf{s}=t_{1}\shortmid\dots\shortmid t_{n}\in V^{*}$ with $n\in\mathbb{N}$, the decoder returns $\texttt{dec}(\mathbf{s})=\texttt{dec}(t_{1})\shortmid\dots\shortmid\texttt{dec}(t_{n})$ using the token definitions. Any string $\bm{\sigma}=c_{1}\shortmid\dots\shortmid c_{m}\in\Sigma^{*}$, $m\in\mathbb{N}$ given to the encoder is tokenized greedily from left to right, each time selecting the token in the vocabulary that represents the most characters starting from the beginning of the string. Specifically, the first token in $\texttt{enc}(\bm{\sigma})$ is the token $t\in V$ such that $\texttt{dec}(t)=c_{1}\shortmid\dots\shortmid c_{i}$, with $i\leq m$, and $\nexists t^{\prime}\in V$ such that $\texttt{dec}(t^{\prime})=c_{1}\shortmid\dots\shortmid c_{j}$ with $i<j\leq m$. In the above selection, if $c_{1}$ is inside a word, then $t$ must contain the special prefix. This process continues in the same manner with the remaining string $c_{i+1}\shortmid\dots\shortmid c_{m}$.
@@ B.2 The Wordpiece tokenization algorithm
Let ${\mathcal{T}}=(\Sigma,V,\texttt{enc},\texttt{dec})$ be a Wordpiece-based tokenizer and $\mathbf{s}\in V^{*}$ a non-canonical token sequence according to ${\mathcal{T}}$. Then, for any token $t\in V$, $\mathbf{s}\shortmid t$ is also non-canonical.
@@
If $\mathbf{s}=t_{1}\shortmid\dots\shortmid t_{n}$ is non-canonical according to Wordpiece, then let $\mathbf{s}^{\prime}=t^{\prime}_{1}\shortmid\dots\shortmid t^{\prime}_{n^{\prime}}$ denote the canonical tokenization of the same character string, $\texttt{dec}(\mathbf{s})=\texttt{dec}(\mathbf{s}^{\prime})$, $n,n^{\prime}\in\mathbb{N}$. Because $\mathbf{s}^{\prime}\neq\mathbf{s}$, there must exist at least one $i\leq\min(n,n^{\prime})$ such that $t_{i}\neq t^{\prime}_{i}$. It is impossible that $t_{i}=t^{\prime}_{i}$ for all $i\in\min(n,n^{\prime})$ but $n\neq n^{\prime}$, because then $\texttt{dec}(\mathbf{s})\neq\texttt{dec}(\mathbf{s}^{\prime})$, as one would be a prefix of the other. Let $t_{i},t^{\prime}_{i}$, with $i\in\min(n,n^{\prime})$ be the first different token between $\mathbf{s}$ and $\mathbf{s}^{\prime}$, i.e., $\forall j<i:t_{j}=t^{\prime}_{j}$ but $t_{i}\neq t^{\prime}_{i}$. Since $\mathbf{s}^{\prime}$ is canonical, it must be that $|t^{\prime}_{i}|>|t_{i}|$, where $|t|=|\texttt{dec}(t)|$ represents the size of token $t$ based on how many characters in $\Sigma$ it encodes. Because $\mathbf{s}$ is a prefix of $\mathbf{s}\shortmid t$, the first $i$ tokens are the same, but $\mathbf{s}\shortmid t$ cannot be canonical because at (token) index $i$ there exists $t^{\prime}_{i}\in V$ that encodes more characters than $t_{i}$, $|t^{\prime}_{i}|>|t_{i}|$. ∎
Appendix C Proof of Theorem 2
@@
Theorem 2. Let $d$ be absolutely continuous with respect to $p$. Moreover, assume that there exist $\mathbf{s}\in V^{*}$ and $t_{1},t_{2}\in V$ such that $\mathbf{s}\shortmid t_{1}$ is non-canonical with $d(\mathbf{s}\shortmid t_{1})>0$ and $\mathbf{s}\shortmid t_{2}$ is canonical with $p(\mathbf{s}\shortmid t_{2})>0$ and $d(\mathbf{s}\shortmid t_{2})>0$. Then, it holds that
@@
Assume there exists $\hat{\mathbf{s}}\in V^{*}$, $t_{1},t_{2}\in V$ such that $\hat{\mathbf{s}}\shortmid t_{1}$ is non-canonical and $d(\hat{\mathbf{s}}\shortmid t_{1})>0$ and $\hat{\mathbf{s}}\shortmid t_{2}$ is canonical and $p(\hat{\mathbf{s}}\shortmid t_{2})>0$ and $d(\hat{\mathbf{s}}\shortmid t_{2})>0$. Given any token sequence $\mathbf{s}\in V^{*}$, let $p_{\mathbf{s}}=P[T|\bm{S}=\mathbf{s}]$ be the true next token distribution and $d_{\mathbf{s}}$, $\tilde{d}_{\mathbf{s}}$ be the next token distribution and canonicalized next token distribution given by the LLM. Then, $d_{\hat{\mathbf{s}}}(t_{1}\,)>0$. Then, we have that $Z=\sum_{t\in V:\ \hat{\mathbf{s}}\shortmid t\text{ is canonical}}d_{\hat{\mathbf{s}}}(t)<1$. By definition of $\tilde{d}_{\hat{\mathbf{s}}}$, this implies that for all $t\in V$ such that $\tilde{d}_{\hat{\mathbf{s}}}(t)>0$, we have that
@@
$\textrm{KL}(p,d)-\textrm{KL}(p,\tilde{d})=$ $\sum_{\mathbf{s}\in V^{*}}p(\mathbf{s})\log\left(\frac{p(\mathbf{s})}{d(\mathbf{s})}\right)-\sum_{\mathbf{s}\in V^{*}}p(\mathbf{s})\log\left(\frac{p(\mathbf{s})}{\tilde{d}(\mathbf{s})}\right)$
$=$ $\sum_{\mathbf{s}\in V^{*}}p(\mathbf{s})\left[\log\left(\frac{p(\mathbf{s})}{d(\mathbf{s})}\right)-\log\left(\frac{p(\mathbf{s})}{\tilde{d}(\mathbf{s})}\right)\right]$
$=$ $\sum_{\mathbf{s}\in V^{*}}p(\mathbf{s})\log\left(\frac{\tilde{d}(\mathbf{s})}{d(\mathbf{s})}\right)$ (6)
$=$ $\sum_{\mathbf{s}\in V^{*}:\ \tilde{d}(\mathbf{s})>0}p(\mathbf{s})\log\left(\frac{\tilde{d}(\mathbf{s})}{d(\mathbf{s})}\right)$ (7)
where the first equations follow from simple manipulations and Eq. 7 follows from the following argument. Whenever $\tilde{d}(\mathbf{s})=0$, it implies that either $d(\mathbf{s})=0$ or that $\mathbf{s}$ is non-canonical. Both cases imply that $p(\mathbf{s})=0$ (either by absolute continuity or non-canonicity). Whenever $p(\mathbf{s})$ and $\tilde{d}(\mathbf{s})$ is zero, the contribution of the corresponding term in Eq. 6 is interpreted as zero because $\lim_{x\to 0^{+}}x\log x=0.$

### v2

@@ B.1 BPE-based tokenizers are non-recovering
    The first $i$ merges in $M_{\mathbf{s}\shortmid t}$ are not the same as in $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$. This means that there exists at least one merge $m\in M_{\mathbf{s}\shortmid t}$ among the first $i$ merges in $M_{\mathbf{s}\shortmid t}$ such that $m\notin[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$. For any such merge $m$, as $m\notin[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, it must hold that $m\in[M_{\mathbf{s}\shortmid t}]_{\bm{t}}$ by Lemma 1. So, in $M_{\mathbf{s}\shortmid t}$, merge $m_{i}$ is preceded by the first i1i-1 merges of $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ and merge $m$. By Lemma 1, $m$ does not affect the tokens that will create $\mathbf{s}$, so the only merges in $M_{\mathbf{s}\shortmid t}$ before $m_{i}$ that affect $\mathbf{s}$ are the first i1i-1 merges of $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, which are the same as $M_{\mathbf{s}^{\prime}}$. Then, as in case 1, merge $m_{i}$ being applied instead of $m^{\prime}_{i}\prec m_{i}$, implies that $M_{\mathbf{s}\shortmid t}$ cannot be the merge sequence that creates the canonical tokenization of $\texttt{dec}(\mathbf{s}\shortmid t)$ according to BPE.
@@
We will now examine the case where $m_{i}=m^{\prime}_{i}$ for all $i\in[\min(n,n^{\prime})]$, $m_{i}\in[M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$, $m^{\prime}_{i}\in M_{\mathbf{s}^{\prime}}$ but $n\neq n^{\prime}$. If $n>n^{\prime}$, then there exists at least one merge that can be applied on $\mathbf{s}^{\prime}$ after all merges of $M_{\mathbf{s}^{\prime}}$ are done, which means that $\mathbf{s}^{\prime}$ cannot be canonical. If $n^{\prime}>n$, then there exists at least one merge that can be applied on $\mathbf{s}$ after all merges of $M_{\mathbf{s}\shortmid t}]_{\mathbf{s}}$ are done. This merge can also be applied on $\mathbf{s}\shortmid t$, which means that $\mathbf{s}\shortmid t$ cannot be canonical.
@@
All cases lead to a contradiction, which implies that $\mathbf{s}\shortmid t$ is non-canonical. We have shown that if $\mathbf{s}$ is non-canonical then $\mathbf{s}\shortmid t$ is also non-canonical. Thus, BPE-based tokenizers are non-recovering. ∎
B.2 Unigram-based tokenizers are non-recovering.
@@
Unigram-based tokenizers are non-recovering.
@@
If $\mathbf{s}$ is non-canonical according to Unigram, then let $\mathbf{s}^{\prime}$ denote the canonical tokenization of the same character string, $\texttt{dec}(\mathbf{s})=\texttt{dec}(\mathbf{s}^{\prime})$. Because $\mathbf{s}^{\prime}$ is canonical, it must be that $r(\mathbf{s}^{\prime})>r(\mathbf{s})$. It follows that $r(\mathbf{s}\shortmid t)=r(\mathbf{s})r(t)<r(\mathbf{s}^{\prime})r(t)=r(\mathbf{s}^{\prime}\shortmid t)$, so $\mathbf{s}\shortmid t$ cannot be the canonical tokenization of $\texttt{dec}(\mathbf{s}\shortmid t)$. We have shown that if $\mathbf{s}$ is non-canonical then $\mathbf{s}\shortmid t$ is also non-canonical. Thus, Unigram-based tokenizers are non-recovering. ∎
B.3 Wordpiece-based tokenizers are non-recovering.
@@
Wordpiece-based tokenizers are non-recovering.
@@
If $\mathbf{s}=t_{1}\shortmid\dots\shortmid t_{n}$ is non-canonical according to Wordpiece, then let $\mathbf{s}^{\prime}=t^{\prime}_{1}\shortmid\dots\shortmid t^{\prime}_{n^{\prime}}$ denote the canonical tokenization of the same character string, $\texttt{dec}(\mathbf{s})=\texttt{dec}(\mathbf{s}^{\prime})$, $n,n^{\prime}\in\mathbb{N}$. Because $\mathbf{s}^{\prime}\neq\mathbf{s}$, there must exist at least one $i\leq\min(n,n^{\prime})$ such that $t_{i}\neq t^{\prime}_{i}$. It is impossible that $t_{i}=t^{\prime}_{i}$ for all $i\in\min(n,n^{\prime})$ but $n\neq n^{\prime}$, because then $\texttt{dec}(\mathbf{s})\neq\texttt{dec}(\mathbf{s}^{\prime})$, as one would be a prefix of the other. Let $t_{i},t^{\prime}_{i}$, with $i\in\min(n,n^{\prime})$ be the first different token between $\mathbf{s}$ and $\mathbf{s}^{\prime}$, i.e., $\forall j<i:t_{j}=t^{\prime}_{j}$ but $t_{i}\neq t^{\prime}_{i}$. Since $\mathbf{s}^{\prime}$ is canonical, it must be that $|t^{\prime}_{i}|>|t_{i}|$, where $|t|=|\texttt{dec}(t)|$ represents the size of token $t$ based on how many characters in $\Sigma$ it encodes. Because $\mathbf{s}$ is a prefix of $\mathbf{s}\shortmid t$, the first $i$ tokens are the same, but $\mathbf{s}\shortmid t$ cannot be canonical because at (token) index $i$ there exists $t^{\prime}_{i}\in V$ that encodes more characters than $t_{i}$, $|t^{\prime}_{i}|>|t_{i}|$. We have shown that if $\mathbf{s}$ is non-canonical then $\mathbf{s}\shortmid t$ is also non-canonical. Thus, Wordpiece-based tokenizers are non-recovering. ∎
Appendix C Non-recoverability under Pretokenization
State-of-the-art LLMs use a tool called pretokenizer in order to split long strings into segments that can be tokenized simultaneously and independent of each other. Formally, the pretokenizer is a function $\texttt{pre}:\Sigma^{+}\rightarrow(\Sigma^{+})^{+}$, where $(\Sigma^{+})^{+}$ represents sequences of strings in $\Sigma^{+}$, such that for string $\bm{\sigma}\in\Sigma^{+}$, $\texttt{pre}(\bm{\sigma})=(\bm{\sigma}_{1},\dots,\bm{\sigma}_{n})$ where $\bm{\sigma}=\bm{\sigma}_{1}\shortmid\dots\shortmid\bm{\sigma}_{n}$, $n\in\mathbb{N}$. The encoder with pretokenizer can then be defined as $\texttt{enc}_{\texttt{pre}}(\bm{\sigma})=\texttt{enc}(\bm{\sigma}_{1})\shortmid\dots\shortmid\texttt{enc}(\bm{\sigma}_{n})$, where $\texttt{enc}:\Sigma^{+}\rightarrow V^{+}$ is an encoder based on the tokenization algorithm used in conjunction with the pretokenizer. We extend the definition of canonical sequences to account for the effect of the pretokenizer.
Definition 8.
Let ${\mathcal{T}}=(\Sigma,V,\texttt{enc}_{\texttt{pre}},\texttt{dec})$ be a tokenizer and $\bm{\sigma}\in\Sigma^{+}$. A tokenization $\mathbf{s}\in V^{+},\texttt{dec}(\mathbf{s})=\bm{\sigma}$ of $\bm{\sigma}$ is canonical if $\mathbf{s}=\texttt{enc}_{\texttt{pre}}(\bm{\sigma})$, where $\texttt{enc}(\bm{\sigma})=\texttt{enc}(\bm{\sigma}_{1})\shortmid\dots\shortmid\texttt{enc}(\bm{\sigma}_{n})$ and $\texttt{pre}(\bm{\sigma})=(\bm{\sigma}_{1},\dots,\bm{\sigma}_{n})$, $n\in\mathbb{N}$.
Pretokenizers typically work by greedily matching prefixes of a string to a regular expression, splitting when the prefix stops matching, and continuing with the remaining suffix. If string $\bm{\sigma}$ is a match, then $\texttt{pre}(\bm{\sigma})=\bm{\sigma}$, and if $\texttt{pre}(\bm{\sigma})=(\bm{\sigma}_{1},\dots,\bm{\sigma}_{n}),n\in\mathbb{N},\bm{\sigma}\in\Sigma^{+}$ then $\texttt{pre}(\bm{\sigma}_{i})=\bm{\sigma}_{i}$ for all $i\in[n]$. Regular expressions used by pretokenizers additionally satisfy a property called closed under prefix (Hayase et al., 2025), though some exceptions apply related to handling whitespace and common english contractions.
Definition 9.
A pretokenizer pre is closed under prefix if for any string $\bm{\sigma}\in\Sigma^{+}$ where $\texttt{pre}(\bm{\sigma})=\bm{\sigma}$ and any prefix $\bm{\sigma}^{\prime}$ of $\bm{\sigma}$, it holds that $\texttt{pre}(\bm{\sigma}^{\prime})=\bm{\sigma}^{\prime}$.
In words, any prefix of a string that is a match to the regular expression is also a match, or equivalently, if a string is not a match then no superstring of it is a match.
We show that tokenizers with pretokenization remain non-recovering when the pretokenizer is closed under prefix.
Theorem 10.
Let ${\mathcal{T}}=(\Sigma,V,\texttt{enc},\texttt{dec})$ be a non-recovering tokenizer and let pre be a pretokenizer closed under prefix. Then, tokenizer ${\mathcal{T}}^{\prime}=(\Sigma,V,\texttt{enc}_{\texttt{pre}},\texttt{dec})$ is non-recovering.
Proof.
Let $\bm{\sigma}=\bm{\sigma}_{1}\shortmid\dots\shortmid\bm{\sigma}_{n}=\texttt{dec}(\mathbf{s})$, where $\texttt{pre}(\bm{\sigma})=(\bm{\sigma}_{1},\dots,\bm{\sigma}_{n})$, $\bm{\sigma}_{i}=\texttt{dec}(\mathbf{s}_{i})$ for all $i\in[n]$ and $\mathbf{s}=\mathbf{s}_{1}\shortmid\dots\shortmid\mathbf{s}_{n}$, $n\in\mathbb{N}$. Additionally, let $\bm{\sigma}_{t}=\texttt{dec}(t)$, so $\bm{\sigma}\shortmid\bm{\sigma}_{t}=\texttt{dec}(\mathbf{s}\shortmid t)$. By definition 9 and because $t$ is a single token, it must hold that either $\texttt{pre}(\bm{\sigma}\shortmid\bm{\sigma}_{t})=(\bm{\sigma}_{1},\dots,\bm{\sigma}_{n},\bm{\sigma}_{t})$ or $\texttt{pre}(\bm{\sigma}\shortmid\bm{\sigma}_{t})=(\bm{\sigma}_{1},\dots,\bm{\sigma}_{n}\shortmid\bm{\sigma}_{t})$. For each $\bm{\sigma}_{i}$, it holds by definition that $\texttt{pre}(\bm{\sigma}_{i})=\bm{\sigma}_{i}$, so $\texttt{enc}_{\texttt{pre}}(\bm{\sigma}_{i})=\texttt{enc}(\texttt{dec}(\mathbf{s}_{i}))$.
By definition 8, for $\mathbf{s}$ to be non-canonical there must be at least one $\mathbf{s}_{i}$ where $\mathbf{s}_{i}\neq\texttt{enc}_{\texttt{pre}}(\texttt{dec}(\mathbf{s}_{i}))=\texttt{enc}(\texttt{dec}(\mathbf{s}_{i}))$, so $\mathbf{s}_{i}$ is non-canonical. As each substring $\bm{\sigma}_{i}$ is tokenized independently, if $i<n$, or if $i=n$ and $\texttt{pre}(\bm{\sigma}\shortmid\bm{\sigma}_{t})=(\bm{\sigma}_{1},\dots,\bm{\sigma}_{n},\bm{\sigma}_{t})$, then $\mathbf{s}_{i}$ is also part of $\mathbf{s}\shortmid t$, so $\mathbf{s}\shortmid t$ is non-canonical. Alternatively, if $i=n$ and $\texttt{pre}(\bm{\sigma}\shortmid\bm{\sigma}_{t})=(\bm{\sigma}_{1},\dots,\bm{\sigma}_{n}\shortmid\bm{\sigma}_{t})$, then because $\mathcal{T}$} is non-recovering, $\mathbf{s}_{n}$ being non-canonical implies that $\mathbf{s}_{n}\shortmid t$ is also non-canonical, therefore $\mathbf{s}\shortmid t$ is non-canonical. In all cases, it holds that if $\mathbf{s}$ is non-canonical then $\mathbf{s}\shortmid t$ is also non-canonical. Thus, tokenizer $\mathcal{T}$$^{\prime}$ is non-recovering. ∎
Appendix D Proof of Theorem 3
@@
Theorem 3. Let $d$ be absolutely continuous with respect to $p$. Moreover, assume that there exist $\mathbf{s}\in V^{+}$ and $t_{1},t_{2}\in V$ such that $\mathbf{s}\shortmid t_{1}$ is non-canonical with $d(\mathbf{s}\shortmid t_{1})>0$ and $\mathbf{s}\shortmid t_{2}$ is canonical with $p(\mathbf{s}\shortmid t_{2})>0$ and $d(\mathbf{s}\shortmid t_{2})>0$. Then, it holds that
@@
Assume there exists $\hat{\mathbf{s}}\in V^{+}$, $t_{1},t_{2}\in V$ such that $\hat{\mathbf{s}}\shortmid t_{1}$ is non-canonical and $d(\hat{\mathbf{s}}\shortmid t_{1})>0$ and $\hat{\mathbf{s}}\shortmid t_{2}$ is canonical and $p(\hat{\mathbf{s}}\shortmid t_{2})>0$ and $d(\hat{\mathbf{s}}\shortmid t_{2})>0$. Given any token sequence $\mathbf{s}\in V^{+}$, let $p_{\mathbf{s}}=P[T|\bm{S}=\mathbf{s}]$ be the true next token distribution and $d_{\mathbf{s}}$, $\tilde{d}_{\mathbf{s}}$ be the next token distribution and canonicalized next token distribution given by the LLM. Then, $d_{\hat{\mathbf{s}}}(t_{1}\,)>0$. Then, we have that $Z=\sum_{t\in V:\ \hat{\mathbf{s}}\shortmid t\text{ is canonical}}d_{\hat{\mathbf{s}}}(t)<1$. By definition of $\tilde{d}_{\hat{\mathbf{s}}}$, this implies that for all $t\in V$ such that $\tilde{d}_{\hat{\mathbf{s}}}(t)>0$, we have that
@@
[…]
$=$ $\sum_{\mathbf{s}\in V^{+}}p(\mathbf{s})\left[\log\left(\frac{p(\mathbf{s})}{d(\mathbf{s})}\right)-\log\left(\frac{p(\mathbf{s})}{\tilde{d}(\mathbf{s})}\right)\right]$
$=$ $\sum_{\mathbf{s}\in V^{+}}p(\mathbf{s})\log\left(\frac{\tilde{d}(\mathbf{s})}{d(\mathbf{s})}\right)$ (6)
$=$ $\sum_{\mathbf{s}\in V^{+}:\ \tilde{d}(\mathbf{s})>0}p(\mathbf{s})\log\left(\frac{\tilde{d}(\mathbf{s})}{d(\mathbf{s})}\right)$ (7)
where the first equations follow from simple manipulations and Eq. 7 follows from the following argument. Whenever $\tilde{d}(\mathbf{s})=0$, it implies that either $d(\mathbf{s})=0$ or that $\mathbf{s}$ is non-canonical. Both cases imply that $p(\mathbf{s})=0$ (either by absolute continuity or non-canonicity). Whenever $p(\mathbf{s})$ and $\tilde{d}(\mathbf{s})$ are zero, the contribution of the corresponding term in Eq. 6 is interpreted as zero because $\lim_{x\to 0^{+}}x\log x=0.$